### v1

@@ 2 Finite element formulation
$a(\bm{u},\bm{v})=\nu(\nabla\bm{u},\nabla\bm{v})+(\bm{a}\cdot\nabla\bm{u},\bm{v}),\quad b(\bm{v},q)=-(q,{\rm div}\bm{v}),\quad\text{with}~{}\bm{u},\bm{v}\in H^{1}(\Omega)^{d},~{}q\in L^{2}(\Omega).$
@@
(1) $\sup_{0\neq\bm{v}_{h}\in\bm{V}_{h}}\frac{b(\bm{v}_{h},p_{h})}{\|\nabla\bm{v}_{h}\|}+s(p_{h},p_{h})^{1/2}\geq\delta_{0}\|p_{h}\|\quad\forall p_{h}\in Q_{h},$
where $\delta_{0}>0$ is a constant that does not depend on the mesh parameter $h$. Suitable definitions of $s(p_{h},q_{h})$ can be found, for example, in [7, 18]. For the analysis that follows, a particular choice of $s(\cdot,\cdot)$ is not important.
@@
$\mathcal{L}(\bm{u}_{h},p_{h};\bm{v}_{h},q_{h})=a(\bm{u}_{h},\bm{v}_{h})+a_{s}(\bm{u}_{h},\bm{v}_{h})+b(\bm{v}_{h},p_{h})+b(\bm{u}_{h},q_{h})-\frac{1}{\nu+\gamma}s(p_{h},q_{h})$
@@
For notation simplicity, let $x=\{\bm{u}_{h},p_{h}\}$ and $y=\{\bm{v}_{h},q_{h}\}$. Based on the symmetric part of $\mathcal{L}(x;y)$, we introduce the following products and norms on $\bm{X}$:
$(x,y)_{\mathcal{L}}=\nu(\nabla\bm{u}_{h},\nabla\bm{v}_{h})+\gamma({\rm div}\bm{u}_{h},{\rm div}\bm{v}_{h})+\frac{1}{\nu+\gamma}\left[(p_{h},q_{h})+s(p_{h},q_{h})\right],\quad\|x\|_{\mathcal{L}}^{2}=(x,x)_{\mathcal{L}}.$
@@
(3) $\|\bm{a}\|_{{L^{\infty}(\Omega)}}=1\quad\text{and}\quad\nu\lesssim 1,~{}~{}\gamma\lesssim 1.$
@@
For $x=\{\bm{u}_{h},p_{h}\},y=\{\bm{v}_{h},q_{h}\}$, we have
$\mathcal{L}(x;y)$ $=a(\bm{u}_{h},\bm{v}_{h})+a_{s}(\bm{u}_{h},\bm{v}_{h})+b(\bm{v}_{h},p_{h})+b(\bm{u}_{h},q_{h})-\tfrac{1}{\nu+\gamma}s(p_{h},q_{h})$
$\leq a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}+\|\bm{a}\|_{L^{\infty}(\Omega)}\|\nabla\bm{u}_{h}\|\|{\bm{v}}_{h}\|+a_{s}(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a_{s}(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}$
$\quad+\|{\rm div}\bm{v}_{h}\|\|p_{h}\|+\|{\rm div}\bm{u}_{h}\|\|q_{h}\|+\frac{1}{\nu+\gamma}s(p_{h},p_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}+\frac{c_{f}}{\nu}\|\sqrt{\nu}\nabla\bm{u}_{h}\|\|\sqrt{\nu}\nabla{\bm{v}}_{h}\|+a_{s}(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a_{s}(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}$
$\quad+(\nu+\gamma)^{-\frac{1}{2}}\|p_{h}\|(\gamma^{\frac{1}{2}}\|{\rm div}\bm{v}_{h}\|+\nu^{\frac{1}{2}}\|\nabla\bm{v}_{h}\|)+(\nu+\gamma)^{-\frac{1}{2}}\|q_{h}\|(\gamma^{\frac{1}{2}}\|{\rm div}\bm{u}_{h}\|$
$\quad+\nu^{\frac{1}{2}}\|\nabla\bm{u}_{h}\|)+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq c_{1}\|\bm{u}_{h},p_{h}\|_{\mathcal{L}}\|\bm{v}_{h},q_{h}\|_{\mathcal{L}},$
@@
We introduce the norm $\|\bm{v}_{h}\|^{2}_{\ast}:=\nu\|\nabla\bm{v}_{h}\|^{2}+\gamma\|{\rm div}\bm{v}_{h}\|^{2}$ and notice that
@@
Thanks to the inf-sup condition (1), for any $p_{h}$, there exists $\bm{w_{h}}$ such that
$b(\bm{w}_{h},p_{h})\geq\left({\delta_{0}}\|p_{h}\|-s(p_{h},p_{h})^{\frac{1}{2}}\right)\|\nabla\bm{w}_{h}\|.$
We may assume that $\|\nabla\bm{w}_{h}\|=\delta_{0}\|p_{h}\|+s(p_{h},p_{h})^{\frac{1}{2}}$. For given $x=\{\bm{u}_{h},p_{h}\}$, let us consider $y=\{\bm{v}_{h},q_{h}\}$ with $\bm{v}_{h}=\bm{u}_{h}+\theta\bm{w}_{h}$ and $q_{h}=-p_{h}$. Then it holds that
$\mathcal{L}$ $(x;y)=a(\bm{u}_{h},\bm{u}_{h}+\theta\bm{w}_{h})+a_{s}(\bm{u}_{h},\bm{u}_{h}+\theta\bm{w}_{h})+b(\bm{u}_{h}+\theta\bm{w}_{h},p_{h})-b(\bm{u}_{h},p_{h})+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})$
$=a(\bm{u}_{h},\bm{u}_{h})+a_{s}(\bm{u}_{h},\bm{u}_{h})+a(\bm{u}_{h},\theta\bm{w}_{h})+a_{s}(\bm{u}_{h},\theta\bm{w}_{h})+b(\theta\bm{w}_{h},p_{h})+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})$
$=\|\bm{u}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})+\nu\theta(\nabla\bm{u}_{h},\nabla\bm{w}_{h})+\gamma\theta({\rm div}\bm{u}_{h},{\rm div}\bm{w}_{h})+\theta(\bm{a}\cdot\nabla\bm{u}_{h},\bm{w}_{h})+b(\theta\bm{w}_{h},p_{h})$
$\geq\|\bm{u}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})-\theta\|\bm{u}_{h}\|_{\ast}\|\bm{w}_{h}\|_{\ast}-\theta\|\nabla\bm{u}_{h}\|\|\bm{w}_{h}\|+\theta\left(\delta_{0}\|p_{h}\|-s(p_{h},p_{h})^{\frac{1}{2}}\right)\|\nabla\bm{w}_{h}\|$
$\geq\|\bm{u}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})-\frac{1}{4}\|\bm{u}_{h}\|_{\ast}^{2}-\theta^{2}\|\bm{w}_{h}\|_{\ast}^{2}-\frac{\nu}{4}\|\nabla\bm{u}_{h}\|^{2}-\frac{\theta^{2}}{\nu}\|\bm{w}_{h}\|^{2}+\theta\left(\delta_{0}^{2}\|p_{h}\|^{2}-s(p_{h},p_{h})\right)$
$\geq\frac{1}{2}\|\bm{u}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(p_{h},p_{h})-\theta^{2}(\nu+\gamma)\|\nabla\bm{w}_{h}\|^{2}-\frac{\theta^{2}}{\nu}c_{f}\|\nabla\bm{w}_{h}\|^{2}+\theta\left(\delta_{0}^{2}\|p_{h}\|^{2}-s(p_{h},p_{h})\right)$
$\geq\frac{1}{2}\|\bm{u}_{h}\|^{2}_{\ast}+(\tfrac{1}{\nu+\gamma}-\theta-2\theta^{2}(\nu+\gamma)-2\theta^{2}\nu^{-1}c_{f})s(p_{h},p_{h})+\theta(1-2\theta(\nu+\gamma)-2\theta\nu^{-1}c_{f})\delta_{0}^{2}\|p_{h}\|^{2}.$
Letting $\theta=\tilde{c}\nu$ with sufficiently small $\tilde{c}\simeq 1$, we obtain
$\mathcal{L}(x;y)\gtrsim\|\bm{u}_{h}\|^{2}_{\ast}+\frac{1}{\nu+\gamma}s(p_{h},p_{h})+\nu\|p_{h}\|^{2}.$
This and $\|\bm{u}_{h},p_{h}\|_{\mathcal{L}}\gtrsim\|\bm{v}_{h},q_{h}\|_{\mathcal{L}}$ prove the result in (5).
@@ 3 Algebraic problem and Schur complement preconditioner
(6) $\mathcal{K}\bm{x}=\begin{pmatrix}A_{\gamma}&~{}B^{T}\\ B&-C\end{pmatrix}\bm{x}=b,$
@@ 3.1 One eigenvalue problem
(11) $Sp=\lambda\widehat{S}p.$
@@
The proof partially follows [38]. For $p\in\mathbb{R}^{m}$, the corresponding finite element function is $p_{h}\in Q_{h}$. For $u\in\mathbb{R}^{n}$, the corresponding finite element function is $\bm{u}_{h}\in\bm{V}_{h}$. According to Bendixson theorem,
$\inf_{0\neq p\in\mathbb{R}^{m}}\frac{\left\langle\frac{1}{2}(S+S^{T})p,p\right\rangle}{\langle\hat{S}p,p\rangle}\leq\Re(\lambda).$
Note that $\left\langle\frac{1}{2}(S+S^{T})p,p\right\rangle=\langle Sp,p\rangle$. For a given $p$, define $u$ by $A_{\gamma}u=B^{T}p$. Then,
$\langle Sp,p\rangle$ $=\langle Bu,p\rangle+\langle Cp,p\rangle,$
$A_{\gamma}u$ $=B^{T}p.$
@@
(12) $\langle Sp,p\rangle$ $=b(\bm{u}_{h},p_{h})+\frac{1}{\nu+\gamma}s(p_{h},p_{h}),$
(13) $b(\bm{v}_{h},p_{h})$ $=a(\bm{u}_{h},\bm{v}_{h})+a_{s}(\bm{u}_{h},\bm{v}_{h}),\quad\forall\bm{v}_{h}\in\bm{V}_{h}.$
Taking $\bm{v}_{h}=\bm{u}_{h}$ in (13) and substituting $b(\bm{u}_{h},p_{h})$ into (12) leads to
$\langle Sp,p\rangle=a(\bm{u}_{h},\bm{u}_{h})+a_{s}(\bm{u}_{h},\bm{u}_{h})+\frac{1}{\nu+\gamma}s(p_{h},p_{h}).$
@@
$\delta\|p_{h}\|^{2}\leq b(\hat{\bm{v}}_{h},p_{h})^{2}+s(p_{h},p_{h}).$
@@
$\delta\|p_{h}\|^{2}$ $\leq(a(\bm{u}_{h},\hat{\bm{v}}_{h})+a_{s}(\bm{u}_{h},\hat{\bm{v}}_{h}))^{2}+s(p_{h},p_{h})$
$\leq\left(a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})^{\frac{1}{2}}+\|\bm{a}_{h}\|_{L^{\infty}(\Omega)}\|\nabla\bm{u}_{h}\|\|\hat{\bm{v}}_{h}\|+\gamma\|{\rm div}\bm{u}\|\|{\rm div}\hat{\bm{v}}_{h}\|\right)^{2}+s(p_{h},p_{h})$
$\leq\left(a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})^{\frac{1}{2}}+c_{f}\|\bm{a}_{h}\|_{L^{\infty}(\Omega)}\|\nabla\bm{u}_{h}\|\|\nabla\hat{\bm{v}}_{h}\|+\gamma\|{\rm div}\bm{u}_{h}\|\|\nabla\hat{\bm{v}}_{h}\|\right)^{2}+s(p_{h},p_{h})$
$\leq\left(\sqrt{\nu}a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}+\frac{c_{f}}{\sqrt{\nu}}a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}+\gamma\|{\rm div}\bm{u}_{h}\|\right)^{2}+s(p_{h},p_{h}),$
@@
$a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})=\nu\|\nabla\hat{\bm{v}}_{h}\|^{2}=\nu,\quad\|{\rm div}\hat{\bm{v}}_{h}\|\leq\|\nabla\hat{\bm{v}}_{h}\|=1,\quad\|\nabla\bm{u}_{h}\|=\frac{1}{\sqrt{\nu}}a(\bm{u}_{h},\bm{u}_{h})^{\frac{1}{2}}.$
@@
$\delta\|p_{h}\|^{2}$ $\leq 2\left((\sqrt{\nu}+c_{f}/\sqrt{\nu})^{2}a(\bm{u}_{h},\bm{u}_{h})+\gamma^{2}\|{\rm div}\bm{u}_{h}\|^{2}\right)+s(p_{h},p_{h})$
$\leq 2\left(2(\nu+c_{f}^{2}\nu^{-1})a(\bm{u}_{h},\bm{u}_{h})+\gamma^{2}a_{s}(\bm{u}_{h},\bm{u}_{h})\right)+(\nu+\gamma)\frac{1}{\nu+\gamma}s(p_{h},p_{h})$
$\leq 4(\nu+1+\gamma^{2}+c_{f}^{2}\nu^{-1})\langle Sp,p\rangle\lesssim(1+\nu^{-1})\langle Sp,p\rangle.$
@@
$\nu(\nu+\gamma)\langle((\nu+\gamma)^{-1}M+C)p,p\rangle\lesssim\langle Sp,p\rangle,$
@@
(14) $\nu(\nu+\gamma)\lesssim\frac{\langle Sp,p\rangle}{\langle\widehat{S}p,p\rangle}.$
@@
$|\lambda|\leq\|\widehat{S}^{-\frac{1}{2}}S\widehat{S}^{-\frac{1}{2}}\|=\sup_{0\neq p,q\in\mathbb{R}^{m}}\frac{\langle\widehat{S}^{-\frac{1}{2}}S\widehat{S}^{-\frac{1}{2}}p,q\rangle}{\|p\|_{\ell^{2}}\|q\|_{\ell^{2}}}=\sup_{0\neq p,q\in\mathbb{R}^{m}}\frac{\langle Sp,q\rangle}{\|\widehat{S}^{\frac{1}{2}}p\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}}.$

### v2

@@ 2 Finite element formulation
$a(\bm{u},\bm{v})=\nu(\nabla\bm{u},\nabla\bm{v})+(\bm{a}\cdot\nabla\bm{u},\bm{v}),\quad b(\bm{v},q)=-(q,{\rm div}\bm{v}),\quad\text{with}~{}\bm{u},\bm{v}\in H^{1}_{0}(\Omega)^{d},~{}q\in L^{2}(\Omega).$
@@
(1) $\sup_{0\neq\bm{v}_{h}\in\bm{V}_{h}}\frac{b(\bm{v}_{h},q_{h})}{\|\nabla\bm{v}_{h}\|}+s(q_{h},q_{h})^{1/2}\geq\delta_{0}\|q_{h}\|\quad\forall q_{h}\in Q_{h},$
where $\delta_{0}>0$ is a constant that does not depend on the mesh parameter $h$. Suitable definitions of $s(g_{h},q_{h})$ can be found, for example, in [7, 18]. For the analysis that follows, a particular choice of $s(\cdot,\cdot)$ is not important.
@@
$\mathcal{L}(\bm{w}_{h},g_{h};\bm{v}_{h},q_{h})=a(\bm{w}_{h},\bm{v}_{h})+a_{s}(\bm{w}_{h},\bm{v}_{h})+b(\bm{v}_{h},g_{h})+b(\bm{w}_{h},q_{h})-\frac{1}{\nu+\gamma}s(g_{h},q_{h})$
@@
For notation simplicity, let $x=\{\bm{w}_{h},g_{h}\}$ and $y=\{\bm{v}_{h},q_{h}\}$. Based on the symmetric part of $\mathcal{L}(x;y)$, we introduce the following products and norms on $\bm{X}$:
$(x,y)_{\mathcal{L}}=\nu(\nabla\bm{w}_{h},\nabla\bm{v}_{h})+\gamma({\rm div}\bm{w}_{h},{\rm div}\bm{v}_{h})+\frac{1}{\nu+\gamma}\left[(g_{h},q_{h})+s(g_{h},q_{h})\right],\quad\|x\|_{\mathcal{L}}^{2}=(x,x)_{\mathcal{L}}.$
@@
(3) $\|\bm{a}\|_{{L^{\infty}(\Omega)}}=1\quad\text{and}\quad\nu\lesssim 1,~{}~{}\gamma\lesssim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\nu^{-1}}.$
@@
For $x=\{\bm{w}_{h},g_{h}\},y=\{\bm{v}_{h},q_{h}\}$, we have
$\mathcal{L}(x;y)$ $=a(\bm{w}_{h},\bm{v}_{h})+a_{s}(\bm{w}_{h},\bm{v}_{h})+b(\bm{v}_{h},g_{h})+b(\bm{w}_{h},q_{h})-\tfrac{1}{\nu+\gamma}s(g_{h},q_{h})$
$\leq a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}+\|\bm{a}\|_{L^{\infty}(\Omega)}\|\nabla\bm{w}_{h}\|\|{\bm{v}}_{h}\|+a_{s}(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a_{s}(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}$
$\quad+\|{\rm div}\bm{v}_{h}\|\|g_{h}\|+\|{\rm div}\bm{w}_{h}\|\|q_{h}\|+\frac{1}{\nu+\gamma}s(g_{h},g_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}+\frac{c_{f}}{\nu}\|\sqrt{\nu}\nabla\bm{w}_{h}\|\|\sqrt{\nu}\nabla{\bm{v}}_{h}\|+a_{s}(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a_{s}(\bm{v}_{h},\bm{v}_{h})^{\frac{1}{2}}$
$\quad+(\nu+\gamma)^{-\frac{1}{2}}\|g_{h}\|(\gamma^{\frac{1}{2}}\|{\rm div}\bm{v}_{h}\|+\nu^{\frac{1}{2}}\|\nabla\bm{v}_{h}\|)+(\nu+\gamma)^{-\frac{1}{2}}\|q_{h}\|(\gamma^{\frac{1}{2}}\|{\rm div}\bm{w}_{h}\|$
$\quad+\nu^{\frac{1}{2}}\|\nabla\bm{w}_{h}\|)+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})^{\frac{1}{2}}s(q_{h},q_{h})^{\frac{1}{2}}$
$\leq c_{1}\|\bm{w}_{h},g_{h}\|_{\mathcal{L}}\|\bm{v}_{h},q_{h}\|_{\mathcal{L}},$
@@
We shall also use $\|\mbox{div}\,{\bm{v}}\|\leq\|\nabla{\bm{v}}\|$ inequality, which follows for $\bm{v}\in H^{1}_{0}(\Omega)^{d}$ from $\Delta=\nabla\mbox{div}-\nabla\times\nabla\times$ identity and an integration by parts argument. We introduce the norm $\|\bm{v}_{h}\|^{2}_{\ast}:=\nu\|\nabla\bm{v}_{h}\|^{2}+\gamma\|{\rm div}\bm{v}_{h}\|^{2}$ and notice that
@@
Thanks to the inf-sup condition (1), for any $g_{h}$, there exists $\hat{\bm{w}}_{h}$ such that
$b(\hat{\bm{w}}_{h},g_{h})\geq\left({\delta_{0}}\|g_{h}\|-s(g_{h},g_{h})^{\frac{1}{2}}\right)\|\nabla\hat{\bm{w}}_{h}\|.$
We may assume that $\|\nabla\hat{\bm{w}}_{h}\|=\delta_{0}\|g_{h}\|+s(g_{h},g_{h})^{\frac{1}{2}}$. For given $x=\{\bm{w}_{h},g_{h}\}$, let us consider $y=\{\bm{v}_{h},q_{h}\}$ with $\bm{v}_{h}=\bm{w}_{h}+\theta\hat{\bm{w}}_{h}$ and $q_{h}=-g_{h}$. Then it holds that
$\mathcal{L}$ $(x;y)=a(\bm{w}_{h},\bm{w}_{h}+\theta\hat{\bm{w}}_{h})+a_{s}(\bm{w}_{h},\bm{w}_{h}+\theta\hat{\bm{w}}_{h})+b(\bm{w}_{h}+\theta\hat{\bm{w}}_{h},g_{h})-b(\bm{w}_{h},g_{h})+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})$
$=a(\bm{w}_{h},\bm{w}_{h})+a_{s}(\bm{w}_{h},\bm{w}_{h})+a(\bm{w}_{h},\theta\hat{\bm{w}}_{h})+a_{s}(\bm{w}_{h},\theta\hat{\bm{w}}_{h})+b(\theta\hat{\bm{w}}_{h},g_{h})+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})$
$=\|\bm{w}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})+\nu\theta(\nabla\bm{w}_{h},\nabla\hat{\bm{w}}_{h})+\gamma\theta({\rm div}\bm{w}_{h},{\rm div}\hat{\bm{w}}_{h})+\theta(\bm{a}\cdot\nabla\bm{w}_{h},\hat{\bm{w}}_{h})+b(\theta\hat{\bm{w}}_{h},g_{h})$
$\geq\|\bm{w}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})-\theta\|\bm{w}_{h}\|_{\ast}\|\hat{\bm{w}}_{h}\|_{\ast}-\theta\|\nabla\bm{w}_{h}\|\|\hat{\bm{w}}_{h}\|+\theta\left(\delta_{0}\|g_{h}\|-s(g_{h},g_{h})^{\frac{1}{2}}\right)\|\nabla\hat{\bm{w}}_{h}\|$
$\geq\|\bm{w}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})-\frac{1}{4}\|\bm{w}_{h}\|_{\ast}^{2}-\theta^{2}\|\hat{\bm{w}}_{h}\|_{\ast}^{2}-\frac{\nu}{4}\|\nabla\bm{w}_{h}\|^{2}-\frac{\theta^{2}}{\nu}\|\hat{\bm{w}}_{h}\|^{2}+\theta\left(\delta_{0}^{2}\|g_{h}\|^{2}-s(g_{h},g_{h})\right)$
$\geq\frac{1}{2}\|\bm{w}_{h}\|^{2}_{\ast}+\tfrac{1}{\nu+\gamma}s(g_{h},g_{h})-\theta^{2}(\nu+\gamma)\|\nabla\hat{\bm{w}}_{h}\|^{2}-\frac{\theta^{2}}{\nu}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{f}^{2}}\|\nabla\hat{\bm{w}}_{h}\|^{2}+\theta\left(\delta_{0}^{2}\|g_{h}\|^{2}-s(g_{h},g_{h})\right)$
$\geq\frac{1}{2}\|\bm{w}_{h}\|^{2}_{\ast}+(\tfrac{1}{\nu+\gamma}-\theta-2\theta^{2}(\nu+\gamma)-2\theta^{2}\nu^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{f}^{2}})s(g_{h},g_{h})+\theta(1-2\theta(\nu+\gamma)-2\theta\nu^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{f}^{2}})\delta_{0}^{2}\|g_{h}\|^{2}.$
Letting $\theta=\tilde{c}\nu$ with sufficiently small $\tilde{c}\simeq 1$ and using (3), we obtain
$\mathcal{L}(x;y)\gtrsim\|\bm{w}_{h}\|^{2}_{\ast}+\frac{1}{\nu+\gamma}s(g_{h},g_{h})+\nu\|g_{h}\|^{2}.$
This and $\|\bm{w}_{h},g_{h}\|_{\mathcal{L}}\gtrsim\|\bm{v}_{h},q_{h}\|_{\mathcal{L}}$ prove the result in (5).
@@ 3 Algebraic problem and Schur complement preconditioner
(6) $\mathcal{K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x}=\begin{pmatrix}A_{\gamma}&~{}B^{T}\\ B&-C\end{pmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x}=b,$
@@ 3.1 One eigenvalue problem
(11) $S\varphi=\lambda\widehat{S}\varphi.$
@@
The proof partially follows [38]. For $\varphi\in\mathbb{R}^{m}$, the corresponding finite element function is $\varphi_{h}\in Q_{h}$. For $w\in\mathbb{R}^{n}$, the corresponding finite element function is $\bm{w}_{h}\in\bm{V}_{h}$. According to Bendixson theorem,
$\inf_{0\neq\varphi\in\mathbb{R}^{m}}\frac{\left\langle\frac{1}{2}(S+S^{T})\varphi,\varphi\right\rangle}{\langle\hat{S}\varphi,\varphi\rangle}\leq\Re(\lambda).$
Note that $\left\langle\frac{1}{2}(S+S^{T})\varphi,\varphi\right\rangle=\langle S\varphi,\varphi\rangle$. For a given $\varphi$, define $w$ by $A_{\gamma}w=B^{T}\varphi$. Then,
$\langle S\varphi,\varphi\rangle$ $=\langle Bw,\varphi\rangle+\langle C\varphi,\varphi\rangle,$
$A_{\gamma}w$ $=B^{T}\varphi.$
@@
(12) $\langle S\varphi,\varphi\rangle$ $=b(\bm{w}_{h},\varphi_{h})+\frac{1}{\nu+\gamma}s(\varphi_{h},\varphi_{h}),$
(13) $b(\bm{v}_{h},\varphi_{h})$ $=a(\bm{w}_{h},\bm{v}_{h})+a_{s}(\bm{w}_{h},\bm{v}_{h}),\quad\forall\bm{v}_{h}\in\bm{V}_{h}.$
Taking $\bm{v}_{h}=\bm{w}_{h}$ in (13) and substituting $b(\bm{w}_{h},\varphi_{h})$ into (12) leads to
$\langle S\varphi,\varphi\rangle=a(\bm{w}_{h},\bm{w}_{h})+a_{s}(\bm{w}_{h},\bm{w}_{h})+\frac{1}{\nu+\gamma}s(\varphi_{h},\varphi_{h}).$
@@
$\delta\|\varphi_{h}\|^{2}\leq b(\hat{\bm{v}}_{h},\varphi_{h})^{2}+s(\varphi_{h},\varphi_{h}).$
@@
$\delta\|\varphi_{h}\|^{2}$ $\leq(a(\bm{w}_{h},\hat{\bm{v}}_{h})+a_{s}(\bm{w}_{h},\hat{\bm{v}}_{h}))^{2}+s(\varphi_{h},\varphi_{h})$
$\leq\left(a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})^{\frac{1}{2}}+\|\bm{a}_{h}\|_{L^{\infty}(\Omega)}\|\nabla\bm{w}_{h}\|\|\hat{\bm{v}}_{h}\|+\gamma\|{\rm div}\bm{w}\|\|{\rm div}\hat{\bm{v}}_{h}\|\right)^{2}+s(\varphi_{h},\varphi_{h})$
$\leq\left(a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})^{\frac{1}{2}}+c_{f}\|\bm{a}_{h}\|_{L^{\infty}(\Omega)}\|\nabla\bm{w}_{h}\|\|\nabla\hat{\bm{v}}_{h}\|+\gamma\|{\rm div}\bm{w}_{h}\|\|\nabla\hat{\bm{v}}_{h}\|\right)^{2}+s(\varphi_{h},\varphi_{h})$
$\leq\left(\sqrt{\nu}a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}+\frac{c_{f}}{\sqrt{\nu}}a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}+\gamma\|{\rm div}\bm{w}_{h}\|\right)^{2}+s(\varphi_{h},\varphi_{h}),$
@@
$a(\hat{\bm{v}}_{h},\hat{\bm{v}}_{h})=\nu\|\nabla\hat{\bm{v}}_{h}\|^{2}=\nu,\quad\|{\rm div}\hat{\bm{v}}_{h}\|\leq\|\nabla\hat{\bm{v}}_{h}\|=1,\quad\|\nabla\bm{w}_{h}\|=\frac{1}{\sqrt{\nu}}a(\bm{w}_{h},\bm{w}_{h})^{\frac{1}{2}}.$
@@
$\delta\|\varphi_{h}\|^{2}$ $\leq 2\left((\sqrt{\nu}+c_{f}/\sqrt{\nu})^{2}a(\bm{w}_{h},\bm{w}_{h})+\gamma^{2}\|{\rm div}\bm{w}_{h}\|^{2}\right)+s(\varphi_{h},\varphi_{h})$
$\leq 2\left(2(\nu+c_{f}^{2}\nu^{-1})a(\bm{w}_{h},\bm{w}_{h})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}a_{s}(\bm{w}_{h},\bm{w}_{h})\right)+(\nu+\gamma)\frac{1}{\nu+\gamma}s(\varphi_{h},\varphi_{h})$
$\leq 4(\nu+1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}+c_{f}^{2}\nu^{-1})\langle S\varphi,\varphi\rangle\lesssim(1+\nu^{-1})\langle S\varphi,\varphi\rangle.$
@@
$\nu(\nu+\gamma)\langle((\nu+\gamma)^{-1}M+C)\varphi,\varphi\rangle\lesssim\langle S\varphi,\varphi\rangle,$
@@
(14) $\nu(\nu+\gamma)\lesssim\frac{\langle S\varphi,\varphi\rangle}{\langle\widehat{S}\varphi,\varphi\rangle}.$
@@
$|\lambda|\leq\|\widehat{S}^{-\frac{1}{2}}S\widehat{S}^{-\frac{1}{2}}\|=\sup_{0\neq\varphi,q\in\mathbb{R}^{m}}\frac{\langle\widehat{S}^{-\frac{1}{2}}S\widehat{S}^{-\frac{1}{2}}\varphi,q\rangle}{\|\varphi\|_{\ell^{2}}\|q\|_{\ell^{2}}}=\sup_{0\neq\varphi,q\in\mathbb{R}^{m}}\frac{\langle S\varphi,q\rangle}{\|\widehat{S}^{\frac{1}{2}}\varphi\|_{\ell^{2}}\|\widehat{S}^{\frac{1}{2}}q\|_{\ell^{2}}}.$